# LeTac-MPC: Learning Model Predictive Control for Tactile-reactive Grasping

### v1

@@ V-C Dynamic Shaking
In this experiment, we have the gripper grasp a box containing beads and shake it violently to explore how LeTac-MPC and baseline methods behave to prevent the box dropping under dynamic shaking. During dynamic shaking, the beads inside also shakes which results in unpredictable state changes in the grasped box. The position trajectory of the end-effector during dynamic shaking is shown in Fig. 13(b). The maximum acceleration can reach to 8.15 m/$\text{s}^{2}$. This makes the task very challenging.
Fig. 14 shows the results of LeTac-MPC. We can see that during the shaking process, the box shook in the hand due to rapid changes in its acceleration and inertia. Correspondingly, the gripper re-grasped based on the tactile feedback and eventually converged. Throughout the entire process, the box did not drop even though it shook in the hand. The focus of the experiment result is on whether the box fall or not. According to Fig. 14, due to the rapid change in the state of the box, minor slip is common, and we are unable to control the slip distance. However, due to the rapid response of LeTac-MPC, it could quickly regrasp the object and converge to a new grasping width to prevent it from falling.
@@
For MPC and PD, we observe that they have similar results, as shown in Fig. 15. We can see that, at the beginning of the experiments, the contact areas are too small compared to the raw images. Therefore, for MPC and PD, the initial converged grasping forces are actually larger than the proper grasping force. We also can get this conclusion by comparing the tactile images of PD and MPC in Fig. 15 with the tactile images of LeTac-MPC in Fig. 14. In the shaking process, the gripper width decreased first due to the sudden change of the grasped object’s state, as seen in Fig. 15. Then, both for PD and MPC, the gripper width increased and converged to a new larger position value, indicating that the initial grasping forces are too large. Since the performace of model-based control methods like PD and MPC rely on the quality of the tactile feedback, their generalization ability is lower than that of LeTac-MPC.
@@
We conducted 10 repetitions of the experiment for each method and calculated the success rate, as shown in Table V. We can see that the success rate of the reactive control methods (LeTac-MPC, MPC, PD) is significantly higher than that of the open-loop grasping. This is because PD and MPC are tend to apply larger force due to the low-quality contact area feedback for the box, as shown in Fig. 15. Additionally, by comparing the tactile images in Figs. 14 and 15, it can be observed that while PD and MPC achieve higher success rates, LeTac-MPC maintains a relatively high success rate while applying significantly less force.
@@
In the daily objects grasping and transportation experiment, we demonstrate that LeTac-MPC’s ability to reactively grasping to a variety of daily objects, since transportation is also a dynamic task. Additionally, this experiment also showcase both PD and MPC exhibit poor generalization across different objects. If a range of diverse objects were introduced in the dynamic shaking experiment, PD and MPC would struggle to grasp some of them effectively, particularly soft ones. On the other hand, open-loop grasping lacks the capability for reactive grasping. Overall, we anticipate that LeTac-MPC will outperform these baseline methods in such scenario. This can also be analogously applied to the obstacle collision experiment.
@@
In Fig. 14, there is an increase to the gripper width. When shaking occurs, the state and in-hand configuration of the object may change. In such cases, the force exerted on GelSight may increase, resulting in changes in the tactile image. Consequently, the controller will increase the gripper width to mitigate this impact. This also could happen in dynamic transportation and obstacle collision experiments. However, the state of the grasped object is constantly changing. If the force exerted on GelSight decreases, the LeTac-MPC will reduce the gripper width to ensure stable grasping. The entire process is dynamic and based on tactile image feedback. Consequently, we can observe that the gripper width sharply decreases and converges to a new position.
@@ V-D Obstacle Collision
In this section, we present the results of the experiment that tests the effect of collisions on LeTac-MPC and baseline methods. In this experiment, we have the robot grasp a screwdriver and then collide it with an obstacle, as shown in Fig. 16. This unexpected collision may occur during some robot manipulation tasks. The ideal grasping controller should be able to resist this type of external impact.
In this experiment, we chose a screwdriver as the grasped object, because it is rigid, regular shape, and lack of surface texture, all of which make it easily being grasped by all three baseline methods (PD, MPC, and open-loop). Otherwise, as demonstrated in the daily objects grasping and transportation experiment, many objects are challenging for PD and MPC to grasp properly and robustly. By using this grasping-feasible object, we further focus on assessing the impact of obstacle collision.
As shown in Fig. 17, in the LeTac-MPC, PD, and MPC experiments, the screwdriver became loose in the hand due to the collision, which results in the gripper rapidly contracting to re-grasp the screwdriver. Depend on the specific grasping and collision situation, reactive grasping controllers re-grasp with corresponding contracting distances and velocities. Therefore, all three methods demonstrates reactive behavior in response to this type of external impacts. On the contrary, as shown in Fig. 18, open-loop grasping cannot handle this type of external impact. When a collision occurs, the grasped object always easily drops. We conducted 10 repetitions of the experiment for each method and calculated the success rate, as shown in Table V. We can see that the success rate of the reactive control methods (LeTac-MPC, MPC, PD) is significantly higher than that of the open-loop grasping.
It is worth noting, as previously mentioned, that the LaTac-MPC exhibits greater fluctuations after convergence compared to PD and MPC. However, these fluctuations are acceptable. As shown in Figs. 14 and 17, the fluctuation range are around ±$\pm 0.05$ mm. As demonstrated in the supplementary video, under such minor fluctuations, changes in the tactile image are virtually indiscernible. Conversely, during dynamic shaking or when the controller reconverges after the grasped object colliding with an external obstacle, the changes in the tactile image become quite visible. The differences between these two indicate that within the range of ±$\pm 0.05$ mm, the state of the grasp remains almost consistent.
Additionally, the LaTac-MPC is more sensitive to changes of the tactile feedback and has a faster response speed. As shown in Fig. 17, at the onset of collision, the tactile features became more pronounced, resulting in stronger feedback. At this point, the LaTac-MPC released the gripper a little bit to maintain a reasonable grasping force. However, with the change of the tactile feedback, it quickly contracted the gripper to stabilize the screwdriver.
@@ VI Discussion and Future work
In this paper, we propose LeTac-MPC, a learning-based model predictive control for tactile-reactive grasping to address the challenges of robotic tactile-reactive grasping for objects with different physical properties and dynamic and force-interactive tasks. The proposed approach features a differentiable MPC layer to model the embeddings extracted by the NN from the tactile feedback, enabling robust reactive grasping of different daily objects. We also present a fully automated data collection pipeline and demonstrate that our trained controller can adapt to various daily objects with different physical properties, sizes, shapes, and surface textures, although the controller is trained with 4 standardized blocks. We perform three tasks: daily objects grasping and transportation, dynamic shaking, and obstacle collision, and compare 4 methods: LeTac-MPC, MPC, PD, and open-loop grasping. Through experimental comparisons, we show that LeTac-MPC has the best performance on dynamic and force-interactive tasks and best generalize ability.
The fundamental idea of our LeTac-MPC lies in learning a unified representation and employing the differentiable MPC layer to utilize it. This approach enables generalization to a variety of daily objects, even when trained on a simplistic dataset. As clearly demonstrated in Fig. 8, the tactile images of each testing object show significant divergence from the training objects. This ability to generalize to unseen objects is a challenge for other learning-based grasping methods, particularly those based on classification, to achieve. It is noteworthy that LeTac-MPC does not explicitly output various physical properties of the grasped object (such as stiffness). Instead, it implicitly extracts these grasp-related properties and then utilizes these implicit features through the MPC layer to generate grasping actions.
@@
We only mount one GelSight sensor on the gripper. This is because a single GelSight sensor suffices to furnish LeTac-MPC with adequate information for effective grasping. Adding a redundant GelSight sensor would increase the computational load for real-time processing, subsequently reducing the control frequency. For the parallel gripper, the GelSight sensors mounted on different fingers can offer detailed information on the shape and texture of an grasped object’s opposing sides, but often provide overlapping information about the information of contact force and states of the grasped object. As we highlight in the paper, LeTac-MPC demonstrates excellent generalization across various shapes and surface textures. Consequently, there is no need for additional modeling, sensing, and extraction of shape and surface texture information. Therefore, LeTac-MPC does not need another GelSight on the gripper.
For slippery object, as long as the object is not extremely slippery, LeTac-MPC can effectively achieve reactive grasping. In our experiments, items such as a box containing beads, a tomato, and an egg, which are all slippery, are successfully addressed by LeTac-MPC. However, this is based on the assumption that the surface of the grasped object is not extremely slippery. In such cases, the tangential force can still cause deformation of the gel surface, which can be captured by the camera in GelSight. If the surface of the object is extremely slippery, and applying sufficient grasping force does not cause enough tangential deformation of the gel surface, the algorithm lacks the necessary information to continue tightening the gripper, leading to a failure in grasping.
It is noteworthy that if we utilize a sensor elastomer with lower stiffness, we anticipate that LeTac-MPC, PD, and MPC will demonstrate enhanced performance in grasping soft objects. They are likely to use less force while achieving more robust grasping, attributed to the greater deformation of the elastomer upon contact with objects. However, a potential issue arises: the excessive softness of the elastomer might cause the grasped object to shake in the hand, which is undesirable for many manipulation tasks. Therefore, selecting an elastomer with the appropriate stiffness is crucial.
@@
1. Firstly, our method cannot perform grasping on very soft objects, such as various types of meat. Due to the much lower stiffness of these objects compared to the GelSight sensor elastomer, the tactile feedback is very subtle and even difficult for human eyes to recognize. To grasp such soft objects, it may be necessary to design new grippers that are more suitable for soft objects and develop sensing and control algorithms that are compatible with new grippers to solve this problem at the system level.
2. Secondly, our algorithm applies to grasping objects with different physical properties, but does not involve specific manipulation operations. In the future, we can explore how to perform complex manipulation tasks on objects with different physical properties based on tactile feedback.
3. In addition, our work only utilizes tactile feedback as the input to the network. However, for more complex manipulation tasks, combining both visual and tactile feedback can provide richer information. In such cases, linear differentiable MPC layer may not have sufficient ability to represent task objectives. To design algorithms that can generalize to various manipulation tasks, incorporating nonlinear optimization layers with both visual and tactile input could be a potential research direction.
4. Vision-based tactile sensors, like GelSight, provide high-dimensional tactile feedback. This high-dimensional feedback contains a wealth of contact information. However, processing such high-dimensional data can challenge the framerate of the entire learning-based control loop, potentially limiting its performance in certain tasks. Tactile sensors based on electrical signals, known for their good sensitivity and wide bandwidth, offer significant advantages. Exploring learning-based reactive grasping combining tactile sensors based on electrical signals and vision is a promising direction.
5. Finally, since the raw image of different GelSight sensors looks slightly different, a model trained with data from one sensor might perform less effectively when transferred to another sensor. However, the imaging principle of GelSight is same and the images from different sensors follow certain patterns and have commonalities. Therefore, training a controller or policy on one sensor and generalizing it to other GelSights presents a promising future research direction.
@@
We thank Yipai Du for helpful discussions about the tactile sensing. We thank Raghava Uppuluri for setting up the initial robot system. This work was partially supported by the United States Department of Agriculture (USDA; grant number 2023-67021-39072). This article solely reflects the opinions and conclusions of its authors and not USDA.

### v2

@@ V-C Dynamic Shaking
In this experiment, we have the gripper grasp a box containing beads and shake it violently to explore how LeTac-MPC and baseline methods behave to prevent the box from falling under dynamic shaking. During dynamic shaking, the beads inside also shake, resulting in unpredictable state changes in the grasped box. The position of the end-effector during dynamic shaking is shown in Fig. 13(b). The maximum acceleration can reach to 8.15 m/$\text{s}^{2}$. This makes the task very challenging.
Fig. 14 shows the results of LeTac-MPC. We can see that during the shaking process, the box shook in the hand due to rapid changes in its acceleration and inertia. Correspondingly, the gripper re-grasped based on the tactile feedback and eventually converged. Throughout the entire process, the box did not drop even though it shook in the hand. The focus of the experiment result is on whether the box falls or not. According to Fig. 14, due to the rapid change in the state of the box, minor slips are common and we cannot control the slip distance. However, due to the rapid response of LeTac-MPC, it could quickly re-grasp the object and converge to a new grasping width to prevent it from falling.
@@
For MPC and PD, we observe that they have similar results, as shown in Fig. 15. We can see that, at the beginning of the experiments, the contact areas are too small compared to the raw images. Therefore, for MPC and PD, the initial convergent grasping forces are actually larger than the proper grasping force. We can also obtain this conclusion by comparing the tactile images of PD and MPC in Fig. 15 with the tactile images of LeTac-MPC in Fig. 14. In the shaking process, the gripper width first decreased due to the sudden change in the state of the grasped object, as seen in Fig. 15. Then, for both PD and MPC, the gripper width increased and converged to a new larger position value, indicating that the initial grasping forces are too large. Since the performance of model-based control methods like PD and MPC rely on the quality of tactile feedback, their generalizability is lower than that of LeTac-MPC.
@@
We conducted 10 repetitions of the experiment for each method and calculated the success rate, as shown in Table V. We can see that the success rate of the reactive control methods (LeTac-MPC, MPC, PD) is significantly higher than that of the open-loop grasping. This is because PD and MPC tend to apply greater force due to the low-quality contact area feedback for the box, as shown in Fig. 15. Additionally, by comparing the tactile images in Figs. 14 and 15, it can be observed that while PD and MPC achieve higher success rates, LeTac-MPC maintains a relatively high success rate while applying significantly less force.
@@
In the daily objects grasping and transportation experiment, we demonstrate that LeTac-MPC’s ability to reactively grasping to a variety of daily objects, since transportation is also a dynamic task. Additionally, this experiment also showcases both PD and MPC exhibit poor generalization across different objects. If a range of diverse objects were introduced in the dynamic shaking experiment, PD and MPC would struggle to grasp some of them effectively, particularly the soft ones. On the other hand, open-loop grasping lacks the capability for reactive grasping. In general, we anticipate that LeTac-MPC will outperform these baseline methods in such scenario. This can also be analogously applied to the obstacle collision experiment.
@@
In Fig. 14, there is an increase to the gripper width. When shaking occurs, the state and in-hand configuration of the object may change. In such cases, the force exerted on GelSight may increase, resulting in changes in the tactile image. Consequently, the controller will increase the gripper width to mitigate this impact. This also could happen in dynamic transportation and obstacle collision experiments. However, the state of the grasped object is constantly changing. If the force exerted on GelSight decreases, the LeTac-MPC will reduce the gripper width to ensure stable grasping. The entire process is dynamic and based on tactile image feedback. Consequently, we can observe that the gripper width is sharply decreased and converges to a new position.
@@ V-D Obstacle Collision
In this section, we present the results of the experiment that tests the effect of collisions on LeTac-MPC and the baseline methods. In this experiment, we have the robot grasp a screwdriver and then collide with an obstacle, as shown in Fig. 16. This unexpected collision may occur during some robot manipulation tasks. The ideal grasping controller should be able to resist this type of external impact.
In this experiment, we chose a screwdriver as the grasped object, because it is rigid, has a regular shape, and lacks surface texture, all of which make it easily grasped by the three baseline methods (PD, MPC, and open-loop). Otherwise, as demonstrated in the daily objects grasping and transportation experiment, many objects are challenging for PD and MPC to grasp properly and robustly. Using this grasp-feasible object, we further focus on assessing the impact of obstacle collision.
As shown in Fig. 17, in the LeTac-MPC, PD, and MPC experiments, the screwdriver became loose in the hand due to the collision, resulting in the gripper rapidly contracting to re-grasp the screwdriver. Depending on the specific grasping and collision situation, reactive grasping controllers re-grasp with corresponding contracting distances and velocities. Therefore, all three methods demonstrate reactive behavior in response to this type of external impacts. On the contrary, as shown in Fig. 18, open-loop grasping cannot handle this type of external impact. When a collision occurs, the grasped object always easily drops. We conducted 10 repetitions of the experiment for each method and calculated the success rate, as shown in Table V. We can see that the success rate of the reactive control methods (LeTac-MPC, MPC, PD) is significantly higher than that of the open-loop grasping.
It is worth noting, as mentioned previously, that LeTac-MPC exhibits higher fluctuations after convergence compared to PD and MPC. However, these fluctuations are acceptable. As shown in Figs. 14 and 17, the fluctuation range is around ±$\pm 0.05$ mm. As demonstrated in the supplementary video, under such minor fluctuations, changes in the tactile image are virtually indiscernible. Conversely, during dynamic shaking or when the controller reconverges after the grasped object collides with an external obstacle, the changes in the tactile image become quite visible. The differences between these two indicate that within the range of ±$\pm 0.05$ mm, the grasp state remains almost constant.
Additionally, the LeTac-MPC is more sensitive to changes in tactile feedback and has a faster response speed. As shown in Fig. 17, at the onset of collision, the tactile features became more pronounced, resulting in stronger feedback. At this point, the LeTac-MPC released the gripper a little bit to maintain a reasonable grasping force. However, with the change of tactile feedback, the gripper quickly contracted to stabilize the screwdriver.
@@ VI Discussion and Future work
In this paper, we propose LeTac-MPC, a learning-based model predictive control for tactile-reactive grasping to address the challenges of robotic tactile-reactive grasping for objects with different physical properties and dynamic and force-interactive tasks. The proposed approach features a differentiable MPC layer to model the embeddings extracted by the NN from the tactile feedback, enabling robust reactive grasping of different daily objects. We also present a fully automated data collection pipeline and demonstrate that our trained controller can adapt to various daily objects with different physical properties, sizes, shapes, and surface textures, although the controller is trained with 4 standardized blocks. We perform three tasks: daily objects grasping and transportation, dynamic shaking, and obstacle collision, and compare 4 methods: LeTac-MPC, MPC, PD, and open-loop grasping. Through experimental comparisons, we show that LeTac-MPC has optimal performance on dynamic and force-interactive tasks and optimal generalizability.
The fundamental idea of our LeTac-MPC lies in learning a unified representation and employing the differentiable MPC layer to utilize it. This approach enables generalization to a variety of daily objects, even when trained on a simplistic dataset. As clearly demonstrated in Fig. 8, the tactile images of each testing object show a significant divergence from the training objects. This ability to generalize to unseen objects is a challenge for other learning-based grasping methods, particularly those based on classification, to achieve. It should be noted that LeTac-MPC does not explicitly output various physical properties of the grasped object (such as stiffness). Instead, it implicitly extracts these grasp-related properties and then utilizes these implicit features through the MPC layer to generate grasping actions.
@@
For slippery object, as long as the object is not extremely slippery, LeTac-MPC can effectively achieve reactive grasping. In our experiments, items such as a box containing beads, a tomato, and an egg, which are all slippery, are successfully addressed by LeTac-MPC. However, this is based on the assumption that the surface of the grasped object is not extremely slippery. In such cases, the tangential force can still cause deformation of the gel surface, which can be captured by the camera in GelSight. If the surface of the object is extremely slippery and the application of sufficient grasping force does not cause enough tangential deformation of the gel surface, the algorithm lacks the necessary information to continue tightening the gripper, leading to a failure of grasping.
It should be noted that if we utilize a sensor elastomer with lower stiffness, we anticipate that LeTac-MPC, PD, and MPC will demonstrate enhanced performance in grasping soft objects. They are likely to use less force while achieving more robust grasping, attributed to the greater deformation of the elastomer upon contact with objects. However, a potential issue arises: the excessive softness of the elastomer might cause the grasped object to shake in the hand, which is undesirable for many manipulation tasks. Therefore, selecting an elastomer with the appropriate stiffness is crucial.
@@
1. First of all, our method cannot grasp very soft objects, such as various types of meat. Due to the much lower stiffness of these objects compared to the GelSight sensor elastomer, the tactile feedback is very subtle and even difficult for human eyes to recognize. To grasp such soft objects, it may be necessary to design new grippers that are more suitable for soft objects and develop sensing and control algorithms that are compatible with new grippers to solve this problem at the system level.
2. Second, our algorithm applies to grasping objects with different physical properties, but does not involve specific manipulation operations. In the future, we can explore how to perform complex manipulation tasks on objects with different physical properties based on tactile feedback.
3. In addition, our work only utilizes tactile feedback as input to the network. However, for more complex manipulation tasks, combining visual and tactile feedback can provide richer information. In such cases, the linear differentiable MPC layer may not have sufficient ability to represent task objectives. To design algorithms that can generalize to various manipulation tasks, incorporating nonlinear optimization layers with visual and tactile input could be a potential research direction.
4. Vision-based tactile sensors, like GelSight, provide high-dimensional tactile feedback. This high-dimensional feedback contains a wealth of contact information. However, processing such high-dimensional data can challenge the frame rate of the entire learning-based control loop, potentially limiting its performance in certain tasks. Tactile sensors based on electrical signals, known for their good sensitivity and wide bandwidth, offer significant advantages. Exploring learning-based reactive grasping that combines tactile sensors based on electrical signals and vision is a promising direction.
5. Finally, since the raw image of different GelSight sensors looks slightly different, a model trained with data from one sensor might perform less effectively when transferred to another sensor. However, the imaging principle of GelSight is the same and the images from different sensors follow certain patterns and have commonalities. Therefore, training a controller or policy on one sensor and generalizing it to other GelSights presents a promising future research direction.
@@
The authors thank Yipai Du for helpful discussions about tactile sensing. The authors thank Raghava Uppuluri for setting up the initial robot system. This work was partially supported by the National Science Foundation (NSF; No. 2423068). This article solely reflects the opinions and conclusions of its authors and not of NSF.